OpenStudy (anonymous):

as a salesperson at trading cards unlimited, justin receives a monthly base pay plus commission on all that his sells. if his sells$500 worth of merchandise in one month, his is paid $410. if his sells $900 of merchandise in one month, his is paid $498. Find justins salary when his sells $3,400 worth of merchandise.

OpenStudy (anonymous):

Total = base pay + commission commission = rate * sales (where rate is the % of sales that Justin gets) At $500: total = base pay + commission on $500 410 = base pay + (rate * 500) At $900: total = base pay + commission on $900 498 = base pay + (rate * 900) We can rewrite the first one as base pay = 410 - (rate * 500) and the second as base pay = 498 - (rate * 900) And since his base pay is always the same, that means 410 - (rate * 500) = 498 - (rate * 900) So (rate * 900) - (rate - 500) = 88 rate * (900 - 500) = 88 rate * 400 = 88 rate = 88 / 400 rate = 0.22, or 22% And we know: 410 = base pay + (rate * 500) so 410 = base pay + (0.22 * 500) 410 = base pay + 110 300 = base pay Now we have our formula in general terms: total = base pay + (rate * sales) total = 300 + (0.22 * sales) So when sales = $3,400: total = 300 + (0.22 * 3400) total = 300 + 748 total = $1,048 Done!

OpenStudy (anonymous):

Tank you !!